OpenStudy (anonymous):

Prove the Pythagorean Theorem using similar triangles. The Pythagorean Theorem states that in a right triangle, the sum of the squares of the lengths of the legs of the triangle equals the squared length of the hypotenuse. Be sure to create and name the appropriate geometric figures.

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

So lets use pythagorean theorem.It says that the sum of a+b must be=c.  so with the pythagorean theorem plug in those numbers and you will get the answer :) btw if you don't know pythagorean theorem it is a^2+b^2=c^2